# Webs Generated by Products of convex and homogeneous Foliations on $\mathbb{P}^{2}$

### v1

@@ Abstract.
This paper investigates flat webs on the projective plane. We present two methods for constructing such webs: the first involves taking the product of two convex reduced foliations and invariant lines, while the second consists of taking the product of finitely many convex homogeneous foliations and invariant lines. In both cases, we demonstrate that the dual web is flat.
@@ Introduction
For a global context, given a $k$-web $\mathcal{W}$ of degree $d$ on the projective plane we can consider the Legendre transform of $\mathcal{W}$, it will be denoted by $\operatorname{Leg}\mathcal{W}$. This is a $d$-web of degree $k$ on the dual projective plane ˇ$\check{\mathbb{P}}^{2}$ which we explain now. Take a generic line $l$ on $\mathbb{P}^{2}$ and consider the tangency locus $\operatorname{Tang}(\mathcal{W},l)=\{p_{1},\cdots,p_{d}\}\subset\mathbb{P}^{2}$ between $\mathcal{W}$ and $l$. We can think the dual $\check{p}_{i}$ as lines on ˇ$\check{\mathbb{P}}^{2}$ passing through the point $l$ of ˇ$\check{\mathbb{P}}^{2}$. Then the set of tangent lines of $\operatorname{Leg}\mathcal{W}$ at $l$ is just $T_{l}\operatorname{Leg}\mathcal{W}=\displaystyle\bigcup_{i=1}^{d}\check{p}_{i}$. More precisely, let $(p,q)$ be the affine chart of ˇ$\check{\mathbb{P}}^{2}$ correspond to the line $\{y=px+q\}\subset\mathbb{P}^{2}$. If $\mathcal{W}$ is defined by an implicit affine equation $F(x,y;p)=0$ with $p=dy/dx$ then $\operatorname{Leg}\mathcal{W}$ is given by the implicit differential equation
@@
Our first result is about the dual of a web generated by product of some invariant lines and two reduced convex foliations under some hypotheses.
@@
Let $\mathcal{F}_{1}$, $\mathcal{F}_{2}$ be reduced convex foliations on $\mathbb{P}_{\mathbb{C}}^{2}$ of degrees $d_{1}$ and $d_{2}$, such that, $\operatorname{Tang}(\mathcal{F}_{1},\mathcal{F}_{2})$ is formed by invariant lines by $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$. If $l_{1},\cdots,l_{n}$ are invariant lines of $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ then the transform $\operatorname{Leg}\mathcal{W}$ of the web $\mathcal{W}=l_{1}\boxtimes\cdots\boxtimes l_{n}\boxtimes\mathcal{F}_{1}\boxtimes\mathcal{F}_{2}$ is flat.
@@
Let $\mathcal{H}_{1},\cdots,\mathcal{H}_{n}$ be convex homogeneous foliations on $\mathbb{P}_{\mathbb{C}}^{2}$ of respective degrees $d_{i}$ with $i=1,\cdots,n$, such that, $\operatorname{Tang}(\mathcal{H}_{i},\mathcal{H}_{j})=L_{\infty}\cup C_{ij}^{inv}$ where $\forall i\neq j$ in $\{1,\cdots,n\}$, $C_{ij}^{inv}$ is a product of invariant lines by $\mathcal{H}_{i}$, $\mathcal{H}_{j}$. If $l_{1},\cdots,l_{k}$ are invariant lines by all the foliations then the dual web $\operatorname{Leg}\mathcal{W}$ of $\mathcal{W}=l_{1}\boxtimes\cdots\boxtimes l_{k}\boxtimes\mathcal{H}_{1}\boxtimes\cdots\boxtimes\mathcal{H}_{n}$ is flat.
These results are expansion of those presented by Bedrouni in [1]. In Theorem A, we added a convex reduced foliation, while the Theorem B can be seen as a generalization of [1, Theorem 2].
@@ 1. Preliminaries
Let $\mathcal{F}$ be a satured foliation on $\mathbb{P}^{2}$, we can define the Gauss map of $\mathcal{F}$ as the rational map $\mathcal{G}_{\mathcal{F}}:\mathbb{P}^{2}\dashrightarrow\check{\mathbb{P}}^{2}$ such that $\mathcal{G}_{\mathcal{F}}(p)=T_{p}\mathcal{F}$, which is well defined outside of $\operatorname{Sing}\mathcal{F}$, called of the singular set of $\mathcal{F}$. If the homogeneous form $\omega=a(x,y,z)dx+b(x,y,z)dy+c(x,y,z)dz$ defines the foliation $\mathcal{F}$, then the Gauss map can be described as
@@
    $I(\mathcal{F})$ does not depend on a particular choice of system of homogeneous coordinates on $\mathbb{P}^{2}$;
@@
    If $C$ is an irreducible algebraic invariant curve of $\mathcal{F}$ then $C\subset I(\mathcal{F})$, if and only if, $C$ is an invariant line;
@@
For a discussion in a more general context and additional details about inflection divisor see [7]. The following result will be useful for us.
@@
We also denote by $\Sigma_{\mathcal{F}}^{\text{rad}}$ the set of radial singularities and by $\check{\Sigma}_{\mathcal{F}}^{\text{rad}}$ the dual lines. It is convenient to define a similar nomenclature: we will denote by $\Sigma_{\mathcal{F}}^{l}$ where $l$ is a line, the set of singularities of $\mathcal{F}$ in $l$ i.e. $\operatorname{Sing}\mathcal{F}\cap l$ and by $\check{\Sigma}_{\mathcal{F}}^{l}$ the respective dual lines. In order to study product of foliations, we shall need the following proposition.
@@
For the our results we will need the following Lemma
@@
Let $\mathcal{F}_{1},\cdots,\mathcal{F}_{k}$ be foliations on $\mathbb{P}_{\mathbb{C}}^{2}$. If $s$ is radial singularity of order $\nu_{i}$ of $\mathcal{F}_{i}$ for $i=1,\cdots,k$, then
$\operatorname{Leg}(\mathcal{F}_{1}\boxtimes\cdots\boxtimes\mathcal{F}_{k})=\mathcal{W}_{1}\boxtimes\mathcal{W}_{2}$
where
@@
    $\mathcal{W}_{1}$ is a $\nu$-web leaving $\check{s}$ invariant, with $\nu=\nu_{1}+\cdots+\nu_{k}$;
@@
    $\check{s}$ is a component totally invariant of $\Delta(\mathcal{W}_{1})$ of minimal order $\nu-1$;
@@
    $\mathcal{W}_{2}$ is transversal to $\check{s}$.
Proof.:  Following [6, Proposition 3.3] we can write, around $s=(0,0)$, the foliations by
$\check{F}_{1}(p,q;x)=q+a_{1}^{1}(p,q)qx+\cdots+a^{1}_{\nu_{1}-1}(p,q)qx^{\nu_{1}-1}+a^{1}_{\nu_{1}}(p,q)x^{\nu_{1}}+\cdots+a^{1}_{d_{1}}(p,q)x^{d_{1}}$
$\vdots$
$\check{F}_{k}(p,q;x)=q+a^{k}_{1}(p,q)qx+\cdots+a^{k}_{\nu_{k}-1}(p,q)qx^{\nu_{k}-1}+a^{k}_{\nu_{k}}(p,q)x^{\nu_{k}}+\cdots+a^{k}_{d_{2}}(p,q)x^{d_{k}}$
with $a^{i}_{\nu_{i}}(p,0)$ not identically zero $\forall i=1,\cdots,k$. Therefore,
$\check{F}(p,q;x)=qG(p,q;x)+a_{\nu}(p,q)x^{\nu}+\cdots+a_{d}(p,q)x^{d}$
where $d=d_{1}+\cdots+d_{k}$ and the degree of $G(p,q;x)$ in $x$ less than or equal to $\nu-1$ and $a_{\nu}(p,0)$ not zero.
Then, we have that $\check{s}$ appears in the discriminant $\Delta(\mathcal{W})$ with minimal order $\nu-1$. Thus, we can to write $\operatorname{Leg}(\mathcal{W})=\mathcal{W}_{1}\boxtimes\mathcal{W}_{2}$ with $\mathcal{W}_{1}$ a $(\nu)$-web leaving $\check{s}$ completely invariant by $\Delta(\mathcal{W}_{1})$ and $\mathcal{W}_{2}$ transversal to $\check{s}$.
@@ 2. Convex Reduced Foliations
Proof.:  As $\mathcal{F}_{1}$, $\mathcal{F}_{2}$ are convex and reduced, we have that all the singularities are non-degenerate ([3, Lemma 6.8]), thus $\Delta(\operatorname{Leg}\mathcal{F}_{i})=\check{\Sigma}_{\mathcal{F}_{i}}^{\text{rad}}$ for each $i=1,2$. furthermore, being $\operatorname{Tang}(\mathcal{F}_{1},\mathcal{F}_{2})=l_{1}\cup\cdots\cup l_{k}$, with each $l_{j}$ invariant, and due to the fact that the Gauss map of invariant lines being points, we obtain the desired equality as a consequence of Proposition 1.2.
@@
Proof.:  By Proposition 2.1, we have that $\Delta(\operatorname{Leg}\mathcal{W})=\check{\Sigma}_{\mathcal{F}_{1}}^{\text{rad}}\cup\check{\Sigma}_{\mathcal{F}_{2}}^{\text{rad}}$. Take $\check{s}\in\check{\Sigma}_{\mathcal{F}_{1}}^{\text{rad}}\setminus\check{\Sigma}_{\mathcal{F}_{2}}^{\text{rad}}$ and a generic line $l$ (particularly, not invariant) passing through $s$. By [6, Proposition 3.3] we can write locally around $l$: $\operatorname{Leg}\mathcal{F}_{1}=\mathcal{W}_{\nu}\boxtimes\mathcal{W}_{d_{1}-\nu}$ with $l$ a totally invariant component of $\Delta(\mathcal{W}_{\nu})$ of minimal multiplicity $\nu-1$ and $\mathcal{W}_{d_{1}-\nu}$ transverse to $\check{s}$. As $s\not\in\Sigma_{\mathcal{F}_{2}}^{\text{rad}}$ we have that $\operatorname{Tang}(\mathcal{F}_{2},l,s)\leq 1$.
  • If $s\not\in\operatorname{Sing}\mathcal{F}_{2}$, we can assume $l\cap\operatorname{Sing}\mathcal{F}_{2}=\emptyset$, then $\operatorname{Leg}\mathcal{F}_{2}$ is completely transversal to $\check{s}$. On the other hand, if one direction of $\mathcal{W}_{d_{1}-\nu}$ and $\operatorname{Leg}\mathcal{F}_{2}$ coincide, this direction would correspond to a point of $\operatorname{Tang}(\mathcal{F}_{1},\mathcal{F}_{2})=l_{1}\cup\cdots\cup l_{k}$ therefore, we would have $l=l_{j}$ for some $j=1,\cdots,k$, which does not occur since $l$ is generic. Thus, $\operatorname{Leg}\mathcal{F}_{2}$ is transverse to $\mathcal{W}_{d-\nu}$ and $\operatorname{Leg}\mathcal{W}=\mathcal{W}_{\nu}\boxtimes(\mathcal{W}_{d_{1}-\nu}\boxtimes\operatorname{Leg}\mathcal{F}_{2})$ with $\mathcal{W}_{d_{1}-\nu}\boxtimes\operatorname{Leg}\mathcal{F}_{2}$ transverse to $\check{s}$. Therefore, by [6, Proposition 2.6], the curvature of $\operatorname{Leg}\mathcal{W}$ is holomorphic along $\check{s}$.
  • If $s\in\operatorname{Sing}\mathcal{F}_{2}$ we have $\operatorname{Tang}(\mathcal{F}_{2},l,s)=1$ and we decompose $\operatorname{Leg}\mathcal{F}_{2}=\mathcal{W}_{1}\boxtimes\mathcal{W}_{d_{2}-1}$. By the previous argument, $\mathcal{W}_{d_{2}-1}$ is transverse to $\mathcal{W}_{d_{1}-\nu}$. Thus, by [2, Proposition 3.9], the curvature of $\operatorname{Leg}\mathcal{W}=(\mathcal{W}_{1}\boxtimes\mathcal{W}_{\nu})\boxtimes(\mathcal{W}_{d_{1}-\nu}\boxtimes\mathcal{W}_{d_{2}-1})$ is holomorphic along $\check{s}$.
It remains to analyze the case $\check{s}\in\check{\Sigma}_{\mathcal{F}_{1}}^{\text{rad}}\cap\check{\Sigma}_{\mathcal{F}_{2}}^{\text{rad}}$, but by Lemma 1.3 and [6, Proposition 2.6], we obtain that $\operatorname{Leg}\mathcal{W}$ is holomorphic along $\check{s}$.
@@
Take now $\mathcal{F}_{l}$, $\mathcal{F}_{d}$ Fermat foliations of degrees $l$ and $d$ with $d=2l-1$. Note that, $\deg(\operatorname{Tang}(\mathcal{F}_{l},\mathcal{F}_{d}))$ $=d+l+1=3l$. On the other hand, with $U_{l-1}(\mathbb{C})$ being the unity group composed by $(l-1)$ roots of unity, we have that $U_{l-1}(\mathbb{C})\subseteq U_{2(l-1)}(\mathbb{C})=U_{d-1}(\mathbb{C})$ thus, $\mathcal{I}_{\mathcal{F}_{l}}\subseteq\mathcal{I}_{\mathcal{F}_{d}}$. Therefore, $\operatorname{Tang}(\mathcal{F}_{l},\mathcal{F}_{d})=\mathcal{I}_{\mathcal{F}_{l}}$ contains only invariant lines of both foliations. We conclude by Proposition 2.2 that $\operatorname{Leg}(\mathcal{F}_{d}\boxtimes\mathcal{F}_{l})$ is flat.

### v2

@@ Abstract.
This paper investigates flat webs on the projective plane. We present two methods for constructing such webs: the first involves taking the product of finitely many convex reduced foliations and invariant lines, while the second consists of taking the product of finitely many convex homogeneous foliations and invariant lines. In both cases, we demonstrate that the dual web is flat.
@@ Introduction
For a global context, given a $k$-web $\mathcal{W}$ of degree $d$ on the projective plane we can consider the Legendre transform of $\mathcal{W}$, it will be denoted by $\operatorname{Leg}\mathcal{W}$. This is a $d$-web of degree $k$ on the dual projective plane ˇ$\check{\mathbb{P}}^{2}$ which we explain now. Take a generic line $l$ on $\mathbb{P}^{2}$ and consider the tangency locus $\operatorname{Tang}(\mathcal{W},l)=\{p_{1},\ldots,p_{d}\}\subset\mathbb{P}^{2}$ between $\mathcal{W}$ and $l$. We can think the dual $\check{p}_{i}$ as lines on ˇ$\check{\mathbb{P}}^{2}$ passing through the point $l$ of ˇ$\check{\mathbb{P}}^{2}$. Then the set of tangent lines of $\operatorname{Leg}\mathcal{W}$ at $l$ is just $T_{l}\operatorname{Leg}\mathcal{W}=\displaystyle\bigcup_{i=1}^{d}\check{p}_{i}$. More precisely, let $(p,q)$ be the affine chart of ˇ$\check{\mathbb{P}}^{2}$ correspond to the line $\{y=px+q\}\subset\mathbb{P}^{2}$. If $\mathcal{W}$ is defined by an implicit affine equation $F(x,y;p)=0$ with $p=dy/dx$ then $\operatorname{Leg}\mathcal{W}$ is given by the implicit differential equation
@@
Our first result is about the dual of a web generated by product of some invariant lines and reduced convex foliations under some hypotheses.
@@
Let $\mathcal{F}_{1},\ldots,\mathcal{F}_{n}$ be reduced convex foliations on $\mathbb{P}^{2}$ of degrees $d_{1},\ldots,d_{n}$ such that $\operatorname{Tang}(\mathcal{F}_{i},\mathcal{F}_{j})$ is formed by invariant lines by $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ where $i\neq j$ in $\{1,\ldots,n\}$. If $l_{1},\ldots,l_{k}$ are invariant lines by all the foliations then the transform $\operatorname{Leg}\mathcal{W}$ of the web $\mathcal{W}=l_{1}\boxtimes\cdots\boxtimes l_{k}\boxtimes\mathcal{F}_{1}\boxtimes\cdots\boxtimes\mathcal{F}_{n}$ is flat.
@@
Let $\mathcal{H}_{1},\ldots,\mathcal{H}_{n}$ be convex homogeneous foliations on $\mathbb{P}^{2}$ of respective degrees $d_{i}$ with $i=1,\ldots,n$, such that, $\operatorname{Tang}(\mathcal{H}_{i},\mathcal{H}_{j})=L_{\infty}\cup C_{ij}^{inv}$ where $\forall i\neq j$ in $\{1,\ldots,n\}$, $C_{ij}^{inv}$ is a product of invariant lines by $\mathcal{H}_{i}$, $\mathcal{H}_{j}$. If $l_{1},\ldots,l_{k}$ are invariant lines by all the foliations then the dual web $\operatorname{Leg}\mathcal{W}$ of $\mathcal{W}=l_{1}\boxtimes\cdots\boxtimes l_{k}\boxtimes\mathcal{H}_{1}\boxtimes\cdots\boxtimes\mathcal{H}_{n}$ is flat.
These results are expansion of those presented by Bedrouni in [1]. The Theorem A can be seen as a generalization of [1, Theorem 1], while the Theorem B is a generalization of [1, Theorem 2]. Furthermore, note that the hypothesis that the tangency locus of two foliations be formed by invariant lines is necessary, as can be seen in the Example 3.6.
@@ 1. Preliminaries
Let $\mathcal{F}$ be a satured foliation on $\mathbb{P}^{2}$, we can define the Gauss map of $\mathcal{F}$ as the rational map $\mathcal{G}_{\mathcal{F}}:\mathbb{P}^{2}\dashrightarrow\check{\mathbb{P}}^{2}$ such that $\mathcal{G}_{\mathcal{F}}(p)=T_{p}\mathcal{F}$, which is well defined outside the singular set of $\mathcal{F}$, $\operatorname{Sing}\mathcal{F}$. If the homogeneous form $\omega=a(x,y,z)dx+b(x,y,z)dy+c(x,y,z)dz$ defines the foliation $\mathcal{F}$, then the Gauss map can be described as
@@
    $I(\mathcal{F})$ does not depend on a particular choice of system of homogeneous coordinates on $\mathbb{P}^{2}$.
@@
    On $\mathbb{P}^{2}\setminus\operatorname{Sing}(\mathcal{F})$, $I(\mathcal{F})$ coincides with the curve described by the inflection points of the leaves of $\mathcal{F}$.
@@
    If $C$ is an irreducible algebraic invariant curve of $\mathcal{F}$ then $C\subset I(\mathcal{F})$, if and only if, $C$ is an invariant line.
  4. (iv)
@@
For a discussion in a more general context and additional details about inflection divisor, see [7]. The following result will be useful for us.
@@
We also denote by $\Sigma_{\mathcal{F}}^{\text{rad}}$ the set of radial singularities and by $\check{\Sigma}_{\mathcal{F}}^{\text{rad}}$ the dual lines. It is convenient to define a similar nomenclature: we will denote by $\Sigma_{\mathcal{F}}^{l}$ where $l$ is a line, the set of singularities of $\mathcal{F}$ in $l$ i.e. $\operatorname{Sing}\mathcal{F}\cap l$ and by $\check{\Sigma}_{\mathcal{F}}^{l}$ the respective dual lines. In order to study the product of foliations, we shall need the following proposition.
@@
For our results, we will need the following lemma.
@@
Let $\mathcal{W}_{1},\mathcal{W}_{2},\ldots\mathcal{W}_{k}$ be $\nu_{i}$-webs and let $C\subset\Delta(\mathcal{W}_{i})$ be an irreducible component totally invariant by $\mathcal{W}_{i}$ and having minimal multiplicity $\nu_{i}-1$ for $i=1,\ldots,k$. Let $\mathcal{W}_{d}$ be a smooth $d$-web transverse to $C$. Then the curvature of $\mathcal{W}=\mathcal{W}_{1}\boxtimes\ldots\boxtimes\mathcal{W}_{k}\boxtimes\mathcal{W}_{d}$ is holomorphic along $C$.
Proof.:  The proof is a simple recurrence of the case $k=2$. So let us assume we are in this case.
Suppose that $1\leq\nu_{1}\leq\nu_{2}$. Let $(U,(z,w))$ be a local coordinate system such that $C\cap U=\{w=0\}$ and
$T\mathcal{W}_{1}|_{U}=\{dw^{\nu_{1}}+w(a_{\nu_{1}-1}(z,w)dw^{\nu_{1}-1}dz+\ldots+a_{0}(z,w)dz^{\nu_{1}})=0\}\hskip 28.45274pt\text{with }w\nmid a_{0}$
$T\mathcal{W}_{2}|_{U}=\{dw^{\nu_{2}}+w(b_{\nu_{2}-1}(z,w)dw^{\nu_{2}-1}dz+\ldots+b_{0}(z,w)dz^{\nu_{2}})=0\}\hskip 28.45274pt\text{with }w\nmid b_{0}$
and
$T\mathcal{W}_{d}|_{U}=\{\prod_{j=1}^{d}(dz+\frac{1}{\nu}g_{j}(z,w)dw)\}$
where $\nu=\nu_{1}\cdot\nu_{2}$. Let $\pi_{i}:\bar{U}\rightarrow U$ be the ramified covering given by $(z,w)=\pi(x,y)=(x,y^{\nu_{i}})$ for $i=1,2$ and $\pi=\pi_{1}\circ\pi_{2}=\pi_{2}\circ\pi_{1}=(x,y^{\nu})$.
The irreducibility of $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ implies that its monodromy group is cyclic and consequently $\pi_{i}^{*}\mathcal{W}_{i}$ is totally decomposable. More precisely, $\pi_{i}^{*}\mathcal{W}_{i}$ is given by
$y^{\nu_{1}(\nu_{1}-2)}dy^{\nu_{1}}+\bar{a}_{\nu_{1}-1}(x,y)y^{(\nu_{1}-1)^{2}}dy^{\nu_{1}-1}dx+\ldots+\bar{a}_{0}dx^{\nu_{1}}=0$
$y^{\nu_{2}(\nu_{2}-2)}dy^{\nu_{2}}+\bar{b}_{\nu_{2}-1}(x,y)y^{(\nu_{2}-1)^{2}}dy^{\nu_{2}-1}dx+\ldots+\bar{b}_{0}dx^{\nu_{2}}=0$
Since $y\nmid\bar{a}_{0}$ and $y\nmid\bar{b}_{0}$ we can write
$\pi_{i}^{*}\mathcal{W}_{i}:\prod_{j=1}^{\nu_{i}}(dx+y^{\nu_{i}-2}f_{i}(x,\xi_{i}^{j}y)\xi_{i}^{-j}dy)=0$
where $\xi_{i}=e^{\frac{2\pi i}{\nu_{i}}}$ and $y\nmid f_{i}$ for each $i=1,2$. Now, we have
$\pi^{*}\mathcal{W}_{1}=\pi_{2}^{*}(\pi_{1}^{*}\mathcal{W}_{1}):\prod_{j=1}^{\nu_{1}}(dx+\nu_{2}y^{\nu-\nu_{2}-1}f_{1}(x,\xi_{1}^{j}y^{\nu 2})\xi_{1}^{-j}dy)=0$
$\pi^{*}\mathcal{W}_{2}=\pi_{1}^{*}(\pi_{2}^{*}\mathcal{W}_{2}):\prod_{j=1}^{\nu_{2}}(dx+\nu_{1}y^{\nu-\nu_{1}-1}f_{2}(x,\xi_{2}^{j}y^{\nu 1})\xi_{2}^{-j}dy)=0$
and
$\pi^{*}\mathcal{W}_{d}:\prod_{j=1}^{d}(dx+y^{\nu-1}g_{j}(x,y^{\nu})dy)=0$
On the other hand, $K(\pi^{*}\mathcal{W})=\sum\limits_{i<j<k}d\eta_{ijk}$ where $\eta_{ijk}$ is the unique $1$-form that satisfies $d(\delta_{rs}\omega_{t})=\eta_{ijk}\wedge\delta_{rs}\omega_{t}$ for each permutation on $(r,s,t)$ of $(i,j,k)$ and the function $\delta_{rs}$ is defined by $\omega_{r}\wedge\omega_{s}=\delta_{rs}(x,y)dx\wedge dy$.
We denote by $\varphi(x,y)=(x,\xi y)$ where $\xi=e^{\frac{2\pi i}{\nu}}$ the deck transformation of $\pi$. We have
$K(\pi^{*}\mathcal{W})=\pi^{*}K(\mathcal{W})=\varphi^{*}\pi^{*}K(\mathcal{W})=\varphi^{*}K(\pi^{*}\mathcal{W}).$
Note that, $\varphi^{*}(y^{n}dx\wedge dy)=y^{n}dx\wedge dy$ if and only if $n\equiv-1\mod\nu$.
If $\sum\eta_{ijk}=Adx+Bdy$ and since $K(\pi^{*}(\mathcal{W}_{2}\boxtimes\mathcal{W}_{d}))$ is holomorphic by [6, Proposition 2.6], we conclude using [6, Lemma 2.2]:
@@
    The order of the poles of $A$ along $\{y=0\}$ is $\leq\nu-\nu_{2}$;
@@
    $B$ is logarithmic along $\{y=0\}$ with constant residue.
Thus we can write $A=\sum\limits_{n=\nu_{2}-\nu}^{\infty}\alpha_{n}(x)y^{n}$ and $B=\sum\limits_{n=-1}^{\infty}\beta_{n}(x)y^{n}$, where $b_{-1}$ is constant. Then $d(\sum\eta_{ijk})=(B_{x}-A_{y})dx\wedge dy$ with polar part given by
$\Omega=\sum\limits_{\nu_{2}-\nu-1}^{-2}(m+1)a_{m+1}y^{m},\hskip 28.45274ptm=n-1$
This fact is summarized with remarks (i) and (ii) which will imply that $d\eta_{ijk}$ is holomorphic along $\{y=0\}$ and consequently $K(\mathcal{W})$ is holomorphic along $C$.
An adaptation of [2, Proposition 2.9] allows us to show a similar result with an additional tangent foliation.
Lemma 1.4.
Let $\mathcal{W}_{1},\ldots,\mathcal{W}_{k}$ be $\nu_{i}$-webs and let $C\subset\Delta(\mathcal{W}_{i})$ for $i=1,\ldots,k$ as in the previous lemma. Let $\mathcal{F}$ be a foliation having $C$ as an invariant curve and let $\mathcal{W}_{d}$ be a smooth $d$-web transverse to $C$. Then the curvature of $\mathcal{W}=\mathcal{F}\boxtimes\mathcal{W}_{1}\boxtimes\ldots\boxtimes\mathcal{W}_{k}\boxtimes\mathcal{W}_{d}$ is holomorphic along $C$.
Some general results about curvature
In order to establish our main results, we shall need the following propositions about some decomposable webs.
Proposition 1.5.
Let $\mathcal{W}=\mathcal{W}_{1}\boxtimes\cdots\boxtimes\mathcal{W}_{n}$ be a product of webs. If for each $i,j,l=1,\ldots,n$ we have:
  1. i.
    $K(\mathcal{W}_{i})=0$;
  2. ii.
    $K(\mathcal{W}_{i}\boxtimes\mathcal{W}_{j})=0$, $i\neq j$;
@@
Then $K(\mathcal{W})=0$.
and
Corollary 1.6.
Let $\mathcal{W}=\mathcal{F}\boxtimes\mathcal{W}_{1}\boxtimes\cdots\boxtimes\mathcal{W}_{n}$ be a product of a foliation and $n$ webs. If for each $i,j,l=1,\ldots,n$ the Proposition 1.5 hypotheses are satisfies and we have:
  1. i.
    $K(\mathcal{F}\boxtimes\mathcal{W}_{i})=0$;
  2. ii.
Then $K(\mathcal{W})=0$.
and
Now a simple computation shows that $K(\mathcal{W})=0$.
Corollary 1.7.
Let $\mathcal{W}=\mathcal{F}_{1}\boxtimes\mathcal{F}_{2}\boxtimes\mathcal{W}_{1}\boxtimes\cdots\boxtimes\mathcal{W}_{n}$ be a product of two foliations and $n$ webs. With the same hypotheses as the previous corollary for $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, and additionally: if $K(\mathcal{F}_{1}\boxtimes\mathcal{F}_{2}\boxtimes\mathcal{W}_{i})=0$ for each $i=1,\ldots,n$. Then $K(\mathcal{W})=0$.
Proof.:  With the same notation as before, we have
Combining this with the information obtained in the previous lemmas, we conclude $K(\mathcal{W})=0$.
Before starting to construct the main results, first note that in the case where we have a foliation and a not invariant line, the dual web will not be flat in general.
@@ 2. Convex Reduced Foliations
Proof.:  As $\mathcal{F}_{1}$, $\mathcal{F}_{2}$ are convex and reduced, we have that all the singularities are non-degenerate ([3, Lemma 6.8]), thus $\Delta(\operatorname{Leg}\mathcal{F}_{i})=\check{\Sigma}_{\mathcal{F}_{i}}^{\text{rad}}$ for each $i=1,2$. furthermore, being $\operatorname{Tang}(\mathcal{F}_{1},\mathcal{F}_{2})=l_{1}\cup\ldots\cup l_{k}$, with each $l_{j}$ invariant, and due to the fact that the Gauss map of invariant lines being points, we obtain the desired equality as a consequence of Proposition 1.2.
@@
Proof.:  By Proposition 2.1, we have that $\Delta(\operatorname{Leg}\mathcal{W})=\check{\Sigma}_{\mathcal{F}_{1}}^{\text{rad}}\cup\check{\Sigma}_{\mathcal{F}_{2}}^{\text{rad}}$. Take $\check{s}\in\check{\Sigma}_{\mathcal{F}_{1}}^{\text{rad}}\setminus\check{\Sigma}_{\mathcal{F}_{2}}^{\text{rad}}$ and a generic line $l$ (particularly, not invariant) passing through $s$. By [6, Proposition 3.3] we can write locally around $l$: $\operatorname{Leg}\mathcal{F}_{1}=\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{d_{1}-\nu_{1}}$ with $l$ a totally invariant component of $\Delta(\mathcal{W}_{\nu_{1}})$ of minimal multiplicity $\nu_{1}-1$ and $\mathcal{W}_{d_{1}-\nu_{1}}$ transverse to $\check{s}$. As $s\not\in\Sigma_{\mathcal{F}_{2}}^{\text{rad}}$ we have that $\operatorname{Tang}(\mathcal{F}_{2},l,s)\leq 1$.
  • If $s\not\in\operatorname{Sing}\mathcal{F}_{2}$, we can assume $l\cap\operatorname{Sing}\mathcal{F}_{2}=\emptyset$, then $\operatorname{Leg}\mathcal{F}_{2}$ is completely transversal to $\check{s}$. On the other hand, if one direction of $\mathcal{W}_{d_{1}-\nu_{1}}$ and $\operatorname{Leg}\mathcal{F}_{2}$ coincide, this direction would correspond to a point of $\operatorname{Tang}(\mathcal{F}_{1},\mathcal{F}_{2})=l_{1}\cup\ldots\cup l_{k}$ therefore, we would have $l=l_{j}$ for some $j=1,\ldots,k$, which does not occur since $l$ is generic. Thus, $\operatorname{Leg}\mathcal{F}_{2}$ is transverse to $\mathcal{W}_{d-\nu_{1}}$ and $\operatorname{Leg}\mathcal{W}=\mathcal{W}_{\nu_{1}}\boxtimes(\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\operatorname{Leg}\mathcal{F}_{2})$ with $\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\operatorname{Leg}\mathcal{F}_{2}$ transverse to $\check{s}$. Therefore, by [6, Proposition 2.6], the curvature of $\operatorname{Leg}\mathcal{W}$ is holomorphic along $\check{s}$.
  • If $s\in\operatorname{Sing}\mathcal{F}_{2}$ we have $\operatorname{Tang}(\mathcal{F}_{2},l,s)=1$ and we decompose $\operatorname{Leg}\mathcal{F}_{2}=\mathcal{W}_{1}\boxtimes\mathcal{W}_{d_{2}-1}$. By the previous argument, $\mathcal{W}_{d_{2}-1}$ is transverse to $\mathcal{W}_{d_{1}-\nu_{1}}$. Thus, by [2, Proposition 3.9], the curvature of $\operatorname{Leg}\mathcal{W}=(\mathcal{W}_{1}\boxtimes\mathcal{W}_{\nu_{1}})\boxtimes(\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\mathcal{W}_{d_{2}-1})$ is holomorphic along $\check{s}$.
Assume now that $s\in\Sigma_{\mathcal{F}_{1}}^{\text{rad}}\cap\Sigma_{\mathcal{F}_{2}}^{\text{rad}}$. In this case we can locally write $\operatorname{Leg}\mathcal{W}=\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{\nu_{2}}\boxtimes\mathcal{W}_{d_{1}+d_{2}-\nu_{1}-\nu_{2}}$ where $\mathcal{W}_{d_{1}+d_{2}-\nu_{1}-\nu_{2}}$ is smooth and transverse to $\check{s}$. We conclude by applying Lemma 1.3 that $K(\operatorname{Leg}\mathcal{W})$ is holomorphic.
@@
Take now $\mathcal{F}_{l}$, $\mathcal{F}_{d}$ Fermat foliations of degrees $l$ and $d$ with $d=2l-1$. Note that, $\deg(\operatorname{Tang}(\mathcal{F}_{l},\mathcal{F}_{d}))$ $=d+l+1=3l$. On the other hand, with $U_{l-1}(\mathbb{C})$ being the unity group composed of $(l-1)$ roots of unity, we have $U_{l-1}(\mathbb{C})\subseteq U_{2(l-1)}(\mathbb{C})=U_{d-1}(\mathbb{C})$ thus, $\mathcal{I}_{\mathcal{F}_{l}}\subseteq\mathcal{I}_{\mathcal{F}_{d}}$. Therefore, $\operatorname{Tang}(\mathcal{F}_{l},\mathcal{F}_{d})=\mathcal{I}_{\mathcal{F}_{l}}$ contains only invariant lines of both foliations. We conclude by Proposition 2.2 that $\operatorname{Leg}(\mathcal{F}_{d}\boxtimes\mathcal{F}_{l})$ is flat.